# Regge trajectories for the hidden bottom tetraquarks in the diquark picture

### v1

@@ III.2 $\lambda-$ and $\rho-$trajectories
Figure 2: The $\lambda-$trajectories for the hidden bottom tetraquarks. Circles represent the predicted data and the blue lines are the $\lambda-$trajectories, see Eq. (II.5). Data are listed in Table 3.
Table 4: Comparison of theoretical predictions for the masses of the hidden bottom tetraquarks (in GeV). $C$ is defined only for $q=q^{\prime}$.
$J^{PC}$ $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ Tetraquark Our FGM Faustov:2021hjs FS Ferretti:2020ewe TR Tiwari:2022azj HK Hadizadeh:2015cvx
@@
When calculating the $\lambda$-mode excitations, the scalar diquark and the scalar antidiquark are used. The radially and orbitally excited states of the $\lambda-$mode are calculated by using Eqs. (II.5), (29) and (30) and parameters in Table 2. The calculated results are listed in Table 3. More results can be easily obtained in a similar manner. The $\lambda-$trajectories for the hidden bottom tetraquarks are shown in Fig. 2.
Table 5: The masses of the $\rho$-mode excited states of the hidden bottom tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $(M^{\prime},\,M^{\prime\prime})$ is for $\left(|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle,\;|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle\right)$. For $(bn)(\bar{b}\bar{n})$ and $(bs)(\bar{b}\bar{s})$, $M^{\prime}=M^{\prime\prime}$. $n=u,\,d$. Eqs. (II.5), (29) and (30) are used.
@@
Figure 3: The $\rho-$trajectories for the hidden bottom tetraquarks. Circles indicate the predicted data and the blue lines are the $\rho-$trajectories, see Eq. (II.5). Data are listed in Table 5.
As discussed in Section C, the values of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ vary with the masses of diquark and antidiquark, $M_{\rho_{1}}$ and $M_{\rho_{2}}$, see Eqs. (II.5), (29) and (30). This means that $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ vary with different states of diquark or antidiquark as we discuss the $\rho-$trajectories, even though $\lambda-$mode remains ground state. Using Eqs. (II.5), (29) and (30) and parameters in Table 2, the masses of the $\rho-$mode excited states are calculated and listed in Table 5.
@@
$\frac{1}{\sqrt{2}}\Big{(}|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$
@@
The obtained results for the $\lambda-$mode excited states are in agreement with other theoretical predictions, see Table 4. To our knowledge, the highly radial and orbital excited states of the $\lambda-$mode states, $|1^{1}s_{0},1^{1}s_{0},4^{1}S_{0}\rangle$, $|1^{1}s_{0},1^{1}s_{0},5^{1}S_{0}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}F_{3}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}G_{4}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}H_{5}\rangle$, and the $\rho-$trajectories of the hidden bottom tetraquarks are discussed for the first time.

### v4

@@ II.4 $\lambda$- and $\rho$-trajectories for the hidden bottom tetraquarks
Figure 2: The $\lambda$-trajectories for the hidden bottom tetraquarks. Circles represent the predicted data and the blue lines are the $\lambda$-trajectories, see Eq. (II.2). Data are listed in Table 2. $M$ is the bound state mass. $N_{r}$ and $L$ are the radial and orbital quantum numbers for the $\lambda$-mode, respectively.
Table 3: Comparison of theoretical predictions for the masses of the hidden bottom tetraquarks (in GeV). The charge conjugation $C$ is defined only for $q=q^{\prime}$.
$J^{PC}$ $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ Tetraquark Our FGS Faustov:2021hjs FS Ferretti:2020ewe TR Tiwari:2022azj HK Hadizadeh:2015cvx
@@
When calculating the $\lambda$-mode excitations, the scalar diquark and the scalar antidiquark are used. The radially and orbitally excited states of the $\lambda$-mode are calculated by using Eqs. (II.2), (42) and (43) and parameters in Table 1. The calculated results are listed in Table 2. More results can be easily obtained in a similar manner. The $\lambda$-trajectories for the hidden bottom tetraquarks are shown in Fig. 2.
Table 4: The masses of the $\rho$-mode excited states of the hidden bottom tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $(M^{\prime},\,M^{\prime\prime})$ is for $\left(|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle,\;|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle\right)$. For $(bn)(\bar{b}\bar{n})$ and $(bs)(\bar{b}\bar{s})$, $M^{\prime}=M^{\prime\prime}$. $n=u,\,d$. Eqs. (II.2), (42) and (43) are used.
@@
Figure 3: The $\rho$-trajectories for the hidden bottom tetraquarks. Circles indicate the predicted data and the blue lines are the $\rho$-trajectories, see Eq. (II.2). Data are listed in Table 4. $M$ is the bound state mass. $n_{r}$ and $l$ are the radial and orbital quantum numbers for the $\rho$-mode, respectively. $1^{1}s_{0}$ and $1^{1}l_{l}$ denote the spin singlet states. $1^{3}s_{1}$ and $1^{3}l_{l+1}$ represent the spin triplet states.
As discussed in Section D, the values of $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ vary with the masses of diquark and antidiquark, $M_{\rho_{1}}$ and $M_{\rho_{2}}$, see Eqs. (II.2), (42) and (43). This means that $c_{fx_{\lambda}}$ and $c_{0x_{\lambda}}$ vary with different states of diquark or antidiquark as we discuss the $\rho$-trajectories, even though $\lambda$-mode remains ground state. Using Eqs. (II.2), (42) and (43) and parameters in Table 1, the masses of the $\rho-$mode excited states are calculated and listed in Table 4.
@@
$|n_{1}^{2s_{1}+1}l_{1j_{1}},$ $n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle^{\prime}$
$=\frac{1}{\sqrt{2}}\Big{(}|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$
@@
The obtained results for the $\lambda$-mode excited states are in agreement with other theoretical predictions, see Table 3. To our knowledge, the highly radially and orbitally excited states of the $\lambda$-mode states, $|1^{1}s_{0},1^{1}s_{0},4^{1}S_{0}\rangle$, $|1^{1}s_{0},1^{1}s_{0},5^{1}S_{0}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}F_{3}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}G_{4}\rangle$, $|1^{1}s_{0},1^{1}s_{0},1^{1}H_{5}\rangle$, and the $\rho$-trajectories of the hidden bottom and charm tetraquarks are discussed for the first time.
The observed $T_{b\bar{b}1}(10610)^{+}$ has been investigated by various theoretical approaches, see Ref. Zhang:2022hfa for references. If $T_{b\bar{b}1}(10610)^{+}$ is a tetraquark, it is the possible candidate for the $|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ state or the $|1^{3}s_{1},1^{3}s_{1},1^{3}S_{1}\rangle$ state, see Table 5. The theoretical predictions agree with the experimental data.
Table 5: Comparison of theoretical predictions and experimental results (in GeV). $C$ is defined only for $q=q^{\prime}$. $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ is defined in Eq. (II.4).
$J^{PC}$ Tetraquark Name State candidate Our PDG ParticleDataGroup:2024cfk FGS Faustov:2021hjs
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$
$1^{+-}$ $(cn)(\bar{c}\bar{n})$ $T_{c\bar{c}1}(3900)^{+}$ $|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ 3.95 3.8871 3.871
$|1^{3}s_{1},1^{3}s_{1},1^{3}S_{1}\rangle$ 3.890
$T_{c\bar{c}1}(4430)^{+}$ $|2^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ 4.44 4.478
$|1^{3}s_{1},2^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ 4.51
$|1^{3}s_{1},1^{1}s_{0},2^{3}S_{1}\rangle^{\prime}$ 4.50 4.431
$|1^{3}s_{1},1^{3}s_{1},2^{3}S_{1}\rangle$ 4.461
$(cn)(\bar{c}\bar{s})$ $T_{c\bar{c}\bar{s}1}(4000)^{+}$ $|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ 4.05 3.988 3.982
$|1^{3}s_{1},1^{3}s_{1},1^{3}S_{1}\rangle$ 4.004
$(bn)(\bar{b}\bar{n})$ $T_{b\bar{b}1}(10610)^{+}$ $|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ 10.49 10.6072 10.492
$|1^{3}s_{1},1^{3}s_{1},1^{3}S_{1}\rangle$ 10.494
II.5 $\lambda$- and $\rho$-trajectories for the hidden charm tetraquarks
Table 6: The masses of the $\lambda$-mode excited states of the hidden charm tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $n=u,\,d$. Eqs. (II.2), (42) and (43) are used.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ $(cn)(\bar{c}\bar{n})$ $(cn)(\bar{c}\bar{s})$ $(cs)(\bar{c}\bar{s})$
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 3.83 3.94 4.05
$|1^{1}s_{0},1^{1}s_{0},2^{1}S_{0}\rangle$ 4.39 4.50 4.60
$|1^{1}s_{0},1^{1}s_{0},3^{1}S_{0}\rangle$ 4.80 4.90 5.00
$|1^{1}s_{0},1^{1}s_{0},4^{1}S_{0}\rangle$ 5.15 5.25 5.34
$|1^{1}s_{0},1^{1}s_{0},5^{1}S_{0}\rangle$ 5.46 5.56 5.65
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 3.84 3.95 4.05
$|1^{1}s_{0},1^{1}s_{0},1^{1}P_{1}\rangle$ 4.23 4.33 4.44
$|1^{1}s_{0},1^{1}s_{0},1^{1}D_{2}\rangle$ 4.52 4.63 4.73
$|1^{1}s_{0},1^{1}s_{0},1^{1}F_{3}\rangle$ 4.78 4.88 4.98
$|1^{1}s_{0},1^{1}s_{0},1^{1}G_{4}\rangle$ 5.01 5.11 5.21
$|1^{1}s_{0},1^{1}s_{0},1^{1}H_{5}\rangle$ 5.23 5.32 5.42
Figure 4: Same as Fig. 2 except for the hidden charm tetraquarks and Table VI.
Table 7: Comparison of theoretical predictions for the masses of the hidden charm tetraquarks (in GeV). $C$ is defined only for $q=q^{\prime}$.
$J^{PC}$ $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ Tetraquark Our FGS Faustov:2021hjs FS Ferretti:2020ewe TR Tiwari:2022azj HK Hadizadeh:2015cvx GLM Giron:2021sla ; Giron:2020fvd
$0^{++}$ $|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ $(cn)(\bar{c}\bar{n})$ 3.83 3.812 3.739 3.8419
$(cn)(\bar{c}\bar{s})$ 3.94 3.922 3.852 3.955 3.9518
$(cs)(\bar{c}\bar{s})$ 4.05 4.051 4.045 3.946
$|1^{1}s_{0},1^{1}s_{0},2^{1}S_{0}\rangle$ $(cn)(\bar{c}\bar{n})$ 4.39 4.375 4.357
$(cn)(\bar{c}\bar{s})$ 4.50 4.481 4.383 4.500
$(cs)(\bar{c}\bar{s})$ 4.60 4.604 4.620 4.558
$1^{--}$ $|1^{1}s_{0},1^{1}s_{0},1^{1}P_{1}\rangle$ $(cn)(\bar{c}\bar{n})$ 4.23 4.244 4.231 4.2244
$(cn)(\bar{c}\bar{s})$ 4.33 4.350 4.234 4.413
$(cs)(\bar{c}\bar{s})$ 4.44 4.466 4.556 4.464
The discussions on the hidden charm tetraquarks are the same as those on the hidden bottom tetraquarks. We assume $m_{u}=m_{d}$, therefore, the masses for diquarks $(cd)$ and $(cu)$ are equal and it holds for the tetraquarks as well.
When considering the $\lambda$-mode excitations, the scalar diquark and the scalar antidiquark are used. Employing Eqs. (II.2), (42), (43), and parameters in Table 1, the masses of the $\lambda$-mode excited states are estimated, see Table 6. The obtained results for the $\lambda$-mode excited states are in agreement with other theoretical predictions, see Table 7. The $\lambda$-trajectories are shown in Fig. 4.
When considering the $\rho$-mode excited states, the $\lambda$-mode state is the ground state, $L=0$ and $N_{r}=0$. The masses of the $\rho$-mode excited states are calculated, see Table 8.
Table 8: The masses of the $\rho$-mode excited states of the hidden charm tetraquarks (in ${{\rm GeV}}$). The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$. $(M^{\prime},\,M^{\prime\prime})$ is for $\big{(}|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle,$ $|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle\big{)}$. For $(cn)(\bar{c}\bar{n})$ and $(cs)(\bar{c}\bar{s})$, $M^{\prime}=M^{\prime\prime}$. $n=u,\,d$. Eqations (II.2), (42), and (43) are used.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ $(cn)(\bar{c}\bar{n})$ $(cn)(\bar{c}\bar{s})$ $(cs)(\bar{c}\bar{s})$
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 3.83 (3.94, 3.94) 4.05
$|2^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 4.40 (4.50, 4.56) 4.67
$|3^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 4.70 (4.80, 4.87) 4.97
$|4^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 4.93 (5.03, 5.11) 5.21
$|5^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 5.12 (5.23, 5.31) 5.41
$|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 3.95 (4.05, 4.04) 4.15
$|2^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 4.44 (4.54, 4.59) 4.69
$|3^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 4.72 (4.83, 4.89) 4.99
$|4^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 4.95 (5.05, 5.12) 5.23
$|5^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 5.14 (5.25, 5.32) 5.42
$|1^{1}s_{0},1^{1}s_{0},1^{1}S_{0}\rangle$ 3.83 (3.94, 3.95) 4.06
$|1^{1}p_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 4.28 (4.38, 4.42) 4.52
$|1^{1}d_{2},1^{1}s_{0},1^{5}S_{2}\rangle$ 4.52 (4.63, 4.66) 4.76
$|1^{1}f_{3},1^{1}s_{0},1^{7}S_{3}\rangle$ 4.71 (4.82, 4.85) 4.95
$|1^{1}g_{4},1^{1}s_{0},1^{9}S_{4}\rangle$ 4.87 (4.98, 5.00) 5.11
$|1^{1}h_{5},1^{1}s_{0},1^{11}S_{5}\rangle$ 5.01 (5.12, 5.14) 5.25
$|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle$ 3.91 (4.02, 4.03) 4.13
$|1^{3}p_{2},1^{1}s_{0},1^{5}S_{2}\rangle$ 4.30 (4.41, 4.44) 4.55
$|1^{3}d_{3},1^{1}s_{0},1^{7}S_{3}\rangle$ 4.54 (4.65, 4.68) 4.78
$|1^{3}f_{4},1^{1}s_{0},1^{9}S_{4}\rangle$ 4.73 (4.83, 4.86) 4.96
$|1^{3}g_{5},1^{1}s_{0},1^{11}S_{5}\rangle$ 4.89 (4.99, 5.02) 5.12
$|1^{3}h_{6},1^{1}s_{0},1^{13}S_{6}\rangle$ 5.03 (5.13, 5.16) 5.26
Figure 5: Same as Fig. 3 except for the hidden charm tetraquarks and Table VIII.
The mass of the state $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ for $(cq)(\bar{c}\bar{q^{\prime}})$ is equal to that of the state $|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle$ for $(cq^{\prime})(\bar{c}\bar{q})$. For the $\rho-$mode excited states, the mass of the state $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$ with definite parity and charge parity (for $q=q^{\prime}$) $1/\sqrt{2}\big{[}(cq)(\bar{c}\bar{q^{\prime}})\pm(cq^{\prime})(\bar{c}\bar{q})\big{]}$ is equal to that of state
$|n_{1}^{2s_{1}+1}l_{1j_{1}},$ $n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle^{\prime}$
$=\frac{1}{\sqrt{2}}\Big{(}|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N^{2j+1}L_{J}\rangle$
±$\pm|n_{2}^{2s_{2}+1}l_{2j_{2}},n_{1}^{2s_{1}+1}l_{1j_{1}},N^{2j+1}L_{J}\rangle\Big{)}$
for $(cq)(\bar{c}\bar{q^{\prime}})$ if $n_{1}^{2s_{1}+1}l_{1j_{1}}$ is not identical to $n_{2}^{2s_{2}+1}l_{2j_{2}}$. The corresponding masses should be $M=(M^{\prime}+M^{\prime\prime})/2$, where $M^{\prime}$ and $M^{\prime\prime}$ are listed in Table 8. Therefore, the corresponding tetraquarks will be degenerate in mass. The $\rho$-trajectories are shown in Fig. 5.
In Ref. Faustov:2021hjs , $T_{c\bar{c}1}(3900)^{+}$ and $T_{c\bar{c}\bar{s}1}(4000)^{+}$ are regarded as $|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ and $|1^{3}s_{1},1^{3}s_{1},1^{3}S_{1}\rangle^{\prime}$ states, respectively. The masses of $T_{c\bar{c}1}(3900)^{+}$ and $T_{c\bar{c}\bar{s}1}(4000)^{+}$ coincide with the theoretical predictions for the $1^{+}$ states, $|1^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ and $|1^{3}s_{1},1^{3}s_{1},1^{3}S_{1}\rangle$, see Table 5. The observed $T_{c\bar{c}1}(4430)^{+}$ has the same $J^{PC}$ and quark content as $T_{c\bar{c}1}(3900)^{+}$. Its mass is $4.478^{+15}_{-18}$ ${\rm GeV}$ ParticleDataGroup:2024cfk . In Ref. Hadizadeh:2015cvx , $T_{c\bar{c}1}(4430)^{+}$ is regarded as the first radially $\lambda$-excited state $|1^{3}s_{1},1^{3}s_{1},2^{3}S_{1}\rangle$ of $(cn)(\bar{c}\bar{n})$. In Ref. Faustov:2021hjs , $T_{c\bar{c}1}(4430)^{+}$ is taken as $|1^{3}s_{1},1^{3}s_{1},2^{3}S_{1}\rangle$ state or $|1^{3}s_{1},1^{1}s_{0},2^{3}S_{1}\rangle^{\prime}$ state. The mass of $T_{c\bar{c}1}(4430)^{+}$ also approximates the mass of the $|2^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ state and the $|1^{3}s_{1},2^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ state, see Table 8. The states $|2^{3}s_{1},1^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ and $|1^{3}s_{1},2^{1}s_{0},1^{3}S_{1}\rangle^{\prime}$ are the $\rho$-excited states. We conclude that $T_{c\bar{c}1}(4430)^{+}$ is possibly the first radial $\lambda$-excitation or the first radial $\rho$-excitation.
II.6 Discussions